OpenStudy (anonymous):

The LCM of 4 and 9 is ____. Numerical Answers Expected!

OpenStudy (anonymous):

@Armyboy

OpenStudy (anonymous):

@fluttershyK

OpenStudy (anonymous):

@emocat1400

OpenStudy (anonymous):

@luckycoins888

OpenStudy (luckycoins888):

Your right Ebay (:

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Which of the following expresses 12 + 42 by using the GCF of the two numbers? 3(9 + 14) 3(4 + 38) 6(2 + 7) 6(6 + 36)

OpenStudy (anonymous):

Which of the following expresses 12 + 42 by using the GCF of the two numbers? 3(9 + 14) 3(4 + 38) 6(2 + 7) 6(6 + 36)

OpenStudy (anonymous):

yes 6(2+7)

OpenStudy (anonymous):

Match the two numbers with their least common multiple (LCM). Match Term Definition 6 and 4 A) 24 6 and 8 B) 30 6 and 10 C) 12

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Which of the sums below can be expressed as 7(4 + 9)? 28 + 9 28 + 63 7 + 63 11 + 16

OpenStudy (anonymous):

hint: 7×(4+9)=7×4+7×9=...?

OpenStudy (anonymous):

we have to apply the distributive property of multiplication over addition

OpenStudy (anonymous):

7*4=...?

OpenStudy (anonymous):

7*9=...?

OpenStudy (fluttershyk):

7(4+9) 28+63

OpenStudy (anonymous):

uhhhh.

OpenStudy (fluttershyk):

@HayHicksxoxo 28+63

OpenStudy (anonymous):

28+63

rebeccaxhawaii (rebeccaxhawaii):

@fluttershyK please refrain from giving out direct answers .. you are very smart and wouldnt wanna see someone like you get suspended. but please explain the question and let the person with the question understand and tell you the answer themselves

OpenStudy (fluttershyk):

@rebeccaxhawaii I di explain and give the answer at the same time on that one I said 7(4+9) 28+36 then he said huhh cuz he didn't relize what I did so I gave the answer again even though I already did

OpenStudy (fluttershyk):

next time you want help on math ask @KamiBug

OpenStudy (fluttershyk):

good bye

OpenStudy (anonymous):

bye